OpenStudy (anonymous):

2/3 of what number is 8?

Parth (parthkohli):

Cool way to ask this: 2/3 multiplied by what is 8?

OpenStudy (phi):

2/3 of what number is 8? I automatically translate "what number" to x I change is to = 2/3 of x = 8 I change "of" to multiply (2/3) x =8

OpenStudy (phi):

once you get \[ \frac{2}{3} x=8\] remember this very useful idea: multiply by the inverse of 2/3 (flip 2/3 to get 3/2) \[ \frac{3}{2}\frac{2}{3} x = \frac{3}{2} \cdot 8\]

Parth (parthkohli):

@amistre64 will answer that question. :)

OpenStudy (amistre64):

@mayankdevnani who needs to get to the solution? you, or the asker? who needs to be able to gain a better understanding of the process and gain confidence in the subject matter, you or the asker?

OpenStudy (amistre64):

let the asker participate in the process and try it out as you guide them, and in the end you can dbl chk their results. That is spose to be the Openstudy way.

OpenStudy (mayankdevnani):

ok @amistre64 thnx for giving me guidelines